OpenStudy (anonymous):

2/(x-y) + 3/(x-y) =1 x+y=2 Solve for X and Y

OpenStudy (saifoo.khan):

5/(x-y) =1 cross multiply. 5 = x-y ================================= Lets solve for x: x = 5 +y Lets solve for y: y = 5 -x

OpenStudy (anonymous):

So 2+3=x-y?

OpenStudy (anonymous):

isnt that what you did

OpenStudy (saifoo.khan):

Oh Right!

OpenStudy (anonymous):

from the first eqn x=5+y put it in the second eqn 5+y+y=2 2y=-3 y=-3/2 put y=-3/2 into the second eqn x-3/2=2 2x-3=4 2x=7 x=7/2 x=7/2 y=-3/2

OpenStudy (anonymous):

how are you getting x=5+y?

OpenStudy (saifoo.khan):

I took the LCM.